# Funny-Valen-Tine: Planning Solution Distribution Enhances Machine Abstract Reasoning Ability

### v2

@@ VI-C Funny for Bongard-Logo problem
This paper suggests that we should also plan the distribution of the pattern representation $\{P_{nm}|m\in[1,M]\}_{v}$ extracted by Valen for visual abstract reasoning problems, so that it is closer to the distribution of the correct solution’s pattern.
However, there are no key or non-key parts in the pattern representation $\{P_{nm}|m\in[1,M]\}_{v}$ encoded by Valen for RPM problems or Bongard-Logo problems, and even currently, the pattern representation $\{P_{nm}|m\in[1,M]\}_{v}$ lacks any interpretability. Therefore, our Funny method is difficult to directly apply to the pattern representations $\{P_{nm}|m\in[1,M]\}_{v}$.
Therefore, we have designed a new method, SBR (Supervised Representation Distribution Planning Method). This method can explicitly model the distribution of $\{P_{nm}|m\in[1,M]\}_{v}$ as a mixture of Gaussian distributions. The idea of SBR is straightforward: SBR defines the distribution form of $\{P_{nm}|m\in[1,M]\}_{v}$ as a mixture of Gaussian distributions and provides additional supervisory signals for the mean and variance of each sub-component in this hypothetical distribution.
When establishing a Gaussian Mixture Model (GMM), providing means, covariance matrices, and component weights is a necessary procedure. However, the objective of this paper is to model the distribution of pattern representations $\{P_{nm}|m\in[1,M]\}_{v}$ as a more interpretable and orderly mixture of Gaussian distributions, which involves altering the original form of the $\{P_{nm}|m\in[1,M]\}_{v}$ distribution. This implies that we need to provide a reasonable number of components for the hypothesized distribution form of $\{P_{nm}|m\in[1,M]\}_{v}$, namely the mixture of Gaussian distributions, assign reasonable means and variances to these components, and provide the covariance matrices between the components. Otherwise, forced and incorrect distribution modeling approaches may detrimentally affect the accuracy of Valen’s reasoning.
It is noteworthy that in the RPM problems, each instance is not only designed with problem statements, option pools, and correct option labels, but also accompanied by metadata. This metadata provides a detailed description of the progressive patterns inherent in each instance. Interestingly, the intended meaning of $\{P_{nm}|m\in[1,M]\}_{v}$ encoded by Valen for an RPM instance also corresponds to a progressive pattern.
Therefore, we decide to use metadata to provide relevant information about the target form of the distribution of $\{P_{nm}|m\in[1,M]\}_{v}$. By enumerating all the metadata within the instances of the RPM problem, the enumerated metadata list can offer ideal information about the component centers for the target distribution form of $\{P_{nm}|m\in[1,M]\}_{v}$. These enumeration results are mutually independent, owing to the characteristics of metadata.
In short, we enumerate all the metadata within the instances of the RPM problem to obtain a metadata list. After processing each element in the list as a Gaussian distribution, the entire list forms a mixture of Gaussian distribution. Subsequently, we naturally assign $\{P_{nm}|m\in[1,M]\}_{v}$ to the corresponding component of this mixture model, thereby achieving our goal of planning the distribution of pattern representations $\{P_{nm}|m\in[1,M]\}_{v}$. This process can be visualized as Figure 19.
@@
In the figure 19, it can be seen that we assume the number of elements in the metadata list to be $S$. Since there are often multiple decoupled progressive patterns within the RPM matrix, such as the “shape” and “line” progressive patterns present in the PGM problem[16], and the related metadata fully describes these progressive patterns, we decide to dissect all the progressive patterns recorded in the metadata and use them to individually guide the assignment of the $M$ progressive patterns stored in the $\{P_{nm}|m\in[1,M]\}_{v}$ to their respective distribution components. In the figure, we dissect the two progressive patterns, “pattern 3” and “pattern S”, recorded in the metadata and use them separately to guide the distribution attribution of $\{P_{n1}|n\in[1,9]\}_{v}$ and $\{P_{n2}|n\in[1,9]\}_{v}$. Evidently, when Valen is combined with the SBR method, the value of $M$ is no longer an arbitrarily set hyperparameter, but rather needs to be determined based on the types of decoupled progressive patterns stored in the metadata. For example, when the SBR method is applied to the PGM problem, the value of $M$ in Valen needs to be set to 2.
Regarding how we process each element in the enumerated metadata list into distributions, we choose to use a standard Transformer-Encoder to encode the elements, which are descriptions of the progressive patterns, into vectors directly. A list containing $S$ elements will be encoded into $S$ vectors, which we denote as $\{q_{s}|s\in[1,S]\}$ here. If the first half of $q_{s}$ is regarded as the mean of a Gaussian distribution and the second half as the logarithm of variance, then $\{q_{s}|s\in[1,S]\}$ can represent a mixture of Gaussian distributions. We can again utilize the reparameterization technique shown in Figure 14 to reparameterize $\{q_{s}|s\in[1,S]\}$ into a new set of vectors, which we denote as $\{Q_{st}|s\in[1,S],t\in[1,T]\}$. In $\{Q_{st}|s\in[1,S],t\in[1,T]\}$, it can be observed that we reparameterized $\{q_{s}|s\in[1,S]\}$ $T$ times, where $T$ is set to 10 in this paper.
@@
Where $\text{meta}_{{m\tilde{s}}}$ represents whether the the $m$-th progressive pattern recorded in the metadata of current instance matches the $\tilde{s}$-th element in the metadata list and $\tau$ is a learnable parameter. As can be seen from the formula (5), considering the overhead of computational resources, we do not individually compress each viewpoint in $\{P_{nm}|n\in[1,9],m\in[1,M]\}_{v}|_{v=1}^{V}$ into the constructed mixture of Gaussian distributions, but instead first calculate the viewpoint average $\{\bar{P}_{nm}|n\in[1,9],m\in[1,M]\}$ of $\{P_{nm}|n\in[1,9],m\in[1,M]\}_{v}$ and then plan the distribution of $\{\bar{P}_{nm}|n\in[1,9],m\in[1,M]\}$.
@@ VIII-A Experiment on RAVEN
Initially, we tested Valen on RAVEN and I-RAVEN datasets to show its potential. For a fair comparison, we used the same settings and equipment as RS-Tran[28] and Triple-CFN[32], including data volume, optimizer parameters, data augmentation, batch size, and other hyperparameters.
TABLE I: Reasoning Accuracies on RAVEN and I-RAVEN.
@@
The experimental records in Table I demonstrate that existing excellent RPM solvers, including Valen, exhibit varying reasoning accuracy on the RAVEN [15] and I-RAVEN [34] datasets. The I-RAVEN dataset was designed to address limitations in the RAVEN’s option pool by enhancing its negative sample setup, such as increasing the number of attributes with shifts and the magnitude of attribute shifts. However, these modifications did not alter the configuration of the primary samples or the distribution of correct solutions. Nonetheless, the same solvers still achieved different reasoning accuracies on RAVEN and I-RAVEN. This observation aligns with our perspective in Tine: the learning objective of probabilistic highlighting solvers is not the distribution of correct solutions, but rather the distribution delineated by both primary and auxiliary samples.
@@ VIII-B Experiment on Bongard-Logo
We conducted experiments on Bongard-Logo. The experimental results are presented in the table II.
TABLE II: Reasoning Accuracies of models on Bongard-Logo.
@@
Model Train FF BA CM NV
MetaOptNet 75.9 60.3 71.6 65.9 67.5
ANIL 69.7 56.6 59.0 59.6 61.0
Meta-Baseline-SC 75.4 66.3 73.3 63.5 63.9
@@
WReN-Bongard 78.7 50.1 50.9 53.8 54.3
SBSD 83.7 75.2 91.5 71.0 74.1
PMoC 92.0 90.6 97.7 77.3 76.0
Triple-CFN 93.2 92.0 98.2 78.0 78.1
@@
Observing table II, The performance of Funny, Valen and Tine were promising.
@@ VIII-C Experiment on PGM
In this paper, we conducted experiments related to Valen on PGM. And the results are recorded in Tabel III.
TABLE III: Reasoning Accuracies of models on PGM.
@@
From the table III, it can be observed that the SBR, Tine, and Funny methods have all contributed to varying degrees of improvement in Valen’s performance on RPM problems. Valen combined with the Funny method is the first to achieve a reasoning accuracy exceeding 99.0% on the PGM problem without relying on any additional supervisory signals. Furthermore, the integration of SBR with Valen is the first approach to utilize metadata as an additional supervisory signal, achieving a reasoning accuracy of 99.4%. Additionally, SBR provides Valen with interpretability by endowing it with a progressive pattern extraction process. The SBR method presented in this paper is not the first to consider utilizing the accompanying metadata within RPM problem instances to assist the solver in reasoning. Previous works[21] have explored various approaches to leveraging metadata, but they ultimately found that forcibly imposing additional metadata learning tasks on their solvers was counterproductive. RS-Tran[28] and Triple-CFN[28] adopted an indirect manner of utilizing metadata. Specifically, RS-Tran utilizes Tranclip[28] to pre-train its representation encoder using metadata, and Triple-CFN required a similar warm-start process. However, the SBR method in this paper is the first to directly utilize metadata to model the distribution of progressive pattern representations, which is an unexplored approach in previous works and has achieved promising results.
We also conducted experiments on the interpolation and extrapolation problems in PGM, and the results are recorded in Table IV. As can be seen from the table IV, the Funny method brings a remarkable improvement to Valen in interpolation.
TABLE IV: Generalization Results of Valen in PGM.
Accuracy(%)
Dataset Valen Valen+Tine Funny+Valen SBR+Valen
Interpolation 81.2 82.8 87.6 92.5
Extrapolation 18.4 18.5 18.5 12.8
@@ IX Conclusion
In conclusion, this paper introduces a novel baseline model, Valen, for solving visual abstract reasoning tasks, particularly Bongard-Logo problems and RPM problems. Valen, a probabilistic-highlighting model, demonstrates remarkable performance in handling both image clustering reasoning and image progression pattern reasoning. The paper further delves into the underlying mechanisms of probability-highlighting solvers, realizing that these solvers approximate solutions to reasoning problem instances as distributions. To bridge the discrepancies between the estimated solution distribution and the true correct solution distribution, the paper introduces three methods: Tine, Funny, and SBR. The Tine method, based on adversarial learning, aims to assist Valen in estimating a solution distribution closer to the correct one. The Funny method models the sample distribution within reasoning problems as a mixture of Gaussian distributions, enabling Valen to more directly capture the true form of the correct solution distribution. The SBR method further extends this idea by modeling the distribution of progressive patterns representation as a mixture of Gaussian distributions. This paper demonstrates the importance of explicitly planning the distribution of solutions in solving visual abstract reasoning problems through the successful application of the SBR, Funny, Valen and Tine.

### v3

@@ VI-D Funny for Bongard-Logo problem
This paper suggests that we should also plan the distribution of the pattern representation $\{P_{nm}|m\in[1,M]\}_{v}$ extracted by Valen for visual abstract reasoning problems, so that it aligns with the distribution of the correct solution’s pattern. However, there are no discernible key or non-key attributes in this pattern representation, and it still lacks sufficient interpretability. Therefore, our Funny method is difficult to directly apply to this planning task
To this end, we have designed a new method, SBR (Supervised Representation Distribution Planning Method). This method can explicitly model the distribution of $\{P_{nm}|m\in[1,M]\}_{v}$ as a mixture of Gaussian distributions[45]. Its idea is straightforward: we reshape the original distribution of $\{P_{nm}|m\in[1,M]\}_{v}$ into an interpretable, well-ordered Gaussian mixture and explicitly supervise the mean and variance of each sub-component in this reshaped distribution.
To establish a mixture of Gaussian distributions, it is necessary to provide means, covariance matrices, and component weights [46]. Consequently, the SBR method needs to specify a reasonable number of components for the distribution of $\{P_{nm}|m\in[1,M]\}_{v}$, assign appropriate means and variances to these components, and specify the covariance structure. Otherwise, any forced or incorrect specification of the reshaped distribution can degrade Valen’s reasoning performance. It is noteworthy that in RPM problems, each instance includes not only the problem statement, option pool, and index of correct option, but also metadata that provide a detailed description of the progressive patterns inherent in the instance. Interestingly, the intended semantics of $\{P_{nm}|m\in[1,M]\}_{v}$ encoded by Valen for an RPM instance also corresponds to these progressive patterns. Therefore, we choose to utilize the metadata to provide detailed information about the target form of the distribution of $\{P_{nm}|m\in[1,M]\}_{v}$.
Specifically, we enumerate all the metadata within RPM instances to obtain a metadata list. After processing each element in the list as a Gaussian distribution, the entire list forms a Gaussian mixture. We then assign the pattern representations $\{P_{nm}|m\in[1,M]\}_{v}$ to the corresponding component of this mixture, thereby achieving our goal of planning the distribution of pattern representations. The assigning process can be visualized as Figure 20.
@@
As shown in this figure, we denote the number of elements in the metadata list to be $S$. Since there are often multiple decoupled progressive patterns within the RPM instance, such as the “shape” and “line” patterns present in the PGM task, and the accompany metadata fully describe these patterns, we decompose all patterns recorded in the metadata and use them individually to guide the assignment of the $M$ patterns stored in the $\{P_{nm}|m\in[1,M]\}_{v}$ to their respective distribution components. Accordingly, we decompose the two progressive patterns, “pattern 3” and “pattern SS”, recorded in the metadata and use them separately to guide the distribution attribution of $\{P_{n1}|n\in[1,9]\}_{v}$ and $\{P_{n2}|n\in[1,9]\}_{v}$. Evidently, when Valen is combined with the SBR method, the value of $M$ is no longer an arbitrary hyperparameter, but must be set to the number of decoupled progressive patterns stored in the metadata; for example, $M=2$ when the SBR method is applied to the PGM task.
Regarding how we process each element in the enumerated metadata list into distributions, we employ a standard Transformer-Encoder to map the elements, which are descriptions of the progressive patterns, into vectors directly. A list of $S$ elements thus yields $S$ vectors, denoted $\{q_{s}|s\in[1,S]\}$. By treating the first half of each $q_{s}$ as the mean and the second half as the logarithm of variance, $\{q_{s}|s\in[1,S]\}$ defines a Gaussian mixture distribution. We again use the reparameterization technique shown in Figure 14 to reparameterize $\{q_{s}|s\in[1,S]\}$ $T$ times into a new set of vectors, denoted as $\{Q_{st}|s\in[1,S],t\in[1,T]\}$, with $T$ is set to 10.
@@
Where $\text{meta}_{{m\tilde{s}}}$ indicates whether the the $m$-th progressive pattern recorded in the metadata of current instance matches the $\tilde{s}-$th element in the metadata list and $\tau$ is a learnable parameter. As shown in Formula (5), to reduce computational overhead we do not compress each viewpoint $v$ of the pattern representations $\{P_{nm}|n\in[1,9],m\in[1,M]\}_{v}|_{v=1}^{V}$ individually into the constructed Gaussian mixture; instead, we compute the viewpoint average $\{\bar{P}_{nm}|n\in[1,9],m\in[1,M]\}$ and plan the distribution of these averaged representations.
@@ VIII-A Experiment on RAVEN
We first evaluated Valen on the RAVEN and I-RAVEN datasets. For a fair comparison, we adopted the same settings and equipment as RS-Tran[32] and Triple-CFN[37], including data volume, optimizer parameters, data augmentation, batch size, and other hyperparameters. For the RAVEN problem, the number of viewpoints $V$ in the Valen model is set to 16, corresponding to a $4\times 4$ patch segmentation.
TABLE I: Reasoning Accuracies Achieved on RAVEN and I-RAVEN.
@@
Human[22] 84.41 95.45 81.82 79.55 86.36 81.81 86.36 81.81
GPT-4V(0-shot)[18] 12.0/- -/- -/- -/- -/- -/- -/- -/-
GPT-4V(1-shot)[18] 12.0/- -/- -/- -/- -/- -/- -/- -/-
Gemini Pro(0-shot)[18] 11.0/- -/- -/- -/- -/- -/- -/- -/-
Gemini Pro(1-shot)[18] 10.0/- -/- -/- -/- -/- -/- -/- -/-
Genome(few-shot)[19] 72.3/- 80.1/- -/- -/- 67.6/- 69.1/- -/- -/-
The experimental records in Table I demonstrate that existing excellent RPM solvers, including Valen, exhibit varying reasoning accuracy on the RAVEN [16] and I-RAVEN [40] datasets. The I-RAVEN dataset was designed to address limitations in the RAVEN’s option pool by enhancing its negative sample setup, such as increasing the number of attributes with shifts and the magnitude of attribute shifts. However, these modifications did not alter the configuration of the primary samples or the distribution of correct solutions. Nonetheless, the same solvers still achieved different reasoning accuracies on RAVEN and I-RAVEN. This observation aligns with our perspective in Tine: the learning objective of probability-highlighting solvers is not the distribution of correct solutions, but rather the distribution delineated by both primary and auxiliary samples.
In Table I, the performance of several LLMs is listed below the human benchmark [22]. Despite high expectations, these models have failed to reach the desired threshold. These findings not only confirm RAVEN’s authority as a benchmark for evaluating AI capability but also demonstrate the superior performance of Valen.
@@ VIII-B Experiment on Bongard-Logo
We conducted experiments on Bongard-Logo. The experiment was conducted in an environment equipped with 4 A100s graphics cards, with a batch size set to 120. The Adam[48] optimizer was chosen, and the initial learning rate was set to $10^{-3}$, with a per-epoch decay method applied for learning rate adjustment, featuring a decay rate of 0.999. It is noteworthy that the Valen model is not sensitive to these parameter settings. The results of the experiment are presented in Table II.
TABLE II: Reasoning Accuracies Achieved on Bongard-Logo.
@@
Model Train $F$ BA CM NV
MetaOptNet[14] 75.9 60.3 71.6 65.9 67.5
ANIL[14] 69.7 56.6 59.0 59.6 61.0
Meta-Baseline-SC[14] 75.4 66.3 73.3 63.5 63.9
@@
WReN-Bongard[14] 78.7 50.1 50.9 53.8 54.3
SBSD[36] 83.7 75.2 91.5 71.0 74.1
PMoC[36] 92.0 90.6 97.7 77.3 76.0
Triple-CFN[37] 93.2 92.0 98.2 78.0 78.1
@@
Human(Expert)[14] - 92.1 99.3 90.7
Human(Amateur)[14] - 88.0 90.0 71.0
As shown in Table II, the performance of Funny, Valen and Tine is promising.
Notably, the transformation that adapts the Bongard Logo problem to Valen’s processing framework (illustrated in Figure 8) is more than merely an RPM-style layout strategy; it is a deliberate redesign grounded in the statistical essence of the Bongard-Logo problem. From a statistical standpoint, the Bongard-Logo problem involves an unknown distribution P(attribute)P({attribute}) to which the attributes of primary group images conform, whereas those of auxiliary group images do not. Solving the Bongard-Logo problem hinges on accurately estimating this latent distribution; the better the estimate, the higher the reasoning accuracy. The Tine method suggests that a probability-highlighting solver is essentially inducing a distribution that assigns high probability to primary group samples and low probability to auxiliary group samples. When this mechanism is applied to Bongard-Logo, its effect is equivalent to estimating P(attribute)P({attribute}), meaning that if such a solver is deployed on this task, it will inherently focus on that core statistical objective. Therefore, the fundamental purpose of our transformation of the Bongard-Logo problem is to externalize the core statistical task of estimating P(attribute)P({attribute}) by leveraging the intrinsic mechanism of probability-highlighting solvers, rather than merely providing a convenient interface.
Behind the seemingly simple rearrangement of the reasoning images lies a methodology that externalizes the statistical essence of reasoning problems. The successful transformation of the Bongard-Logo task not only yields a new approach to the task itself but also offers guidance for other reasoning problems whose solutions hinge on estimating attribute distributions.
@@ VIII-C Experiment on PGM
In this paper, we conducted experiments related to Valen on PGM. For the PGM problem, the number of viewpoints $V$ in the Valen model is set to 16, corresponding to a $4\times 4$ patch division. For a fair comparison, this paper adopts the same configuration as RS-Tran[32] and Triple-CFN[37] in terms of hardware selection, data volume, optimizer parameters, data augmentation methods, batch size, and other settings. And the results are recorded in Table III.
TABLE III: Reasoning Accuracies Achieved on PGM.
@@
Human[17] >>80.0
From Table III, it can be observed that the SBR, Tine, and Funny methods have all contributed to varying degrees of improvement in Valen’s performance on RPM problems. The combination of Valen and the Funny method has achieved a reasoning accuracy exceeding 99.3% on the PGM problem for the first time, without relying on any additional supervisory signals. Furthermore, the combination of SBR with Valen represents the first method that achieves a 99.4% reasoning accuracy by utilizing metadata as an additional supervisory signal. The SBR method presented in this paper is not the first to consider utilizing the accompanying metadata within RPM problem instances to assist the solver in reasoning. Previous works[25, 32, 37] have explored various approaches to leveraging metadata, but they ultimately found that directly and forcibly imposing additional metadata learning tasks on their solvers was counterproductive[25]. RS-Tran[32] and Triple-CFN[37] adopted an indirect manner of utilizing metadata. Specifically, RS-Tran utilizes Tranclip[32] to pre-train its representation encoder using metadata, and Triple-CFN required a similar warm-start process. However, the SBR method in this paper is the first to directly utilize metadata to model the distribution of progressive pattern representations, which is an unexplored approach in previous works and has achieved remarkable results.
We also carried out experiments on the generalization issue of PGM, and the findings are documented in Table IV. As evident from Table IV, the Funny method significantly enhances Valen’s performance on the attribute generalization task Interpolation, whereas SBR exhibits notable advancements in the two progressive generalization patterns, Held-out Pairs of Triples and Held-out Attribute Pairs.
TABLE IV: Reasoning Accuracies Achieved on PGM Generalization Tasks.
Dataset
Model/
Task
Interpolation Extrapolation
Held-out
Attribute
shape-color
Held-out
Attribute
line-type
Held-out
Triples
Held-out
Pairs of
Triples
Held-out
Attribute
Pairs
MRNet[25]
68.1 19.2 16.9 30.1 25.9 55.3 38.4
RS-Tran[32]
77.2 19.2 12.9 24.7 22.2 43.6 28.4
Triple-CFN+Re-space layer[37]
80.4 18.4 12.6 25.2 22.0 44.5 29.2
Valen 81.2 18.4 12.8 26.0 22.8 44.8 29.5
Valen+Tine
82.8 18.5 13.2 26.4 23.3 45.2 30.0
Funny+Valen
87.6 18.5 13.6 27.2 24.1 45.2 30.2
SBR+Valen
92.5 12.8 13.8 26.8 28.2 98.1 98.2
Table IV also indicates that MRNet demonstrates a slight advantage in the remaining generalization tasks; however, this paper takes a cautious stance, arguing that it is difficult to unequivocally determine which solver is better when the reasoning accuracy hovers around 30%, as a reasoning accuracy below this threshold already underscores the limitations of the solvers, making it inadequate to discuss superiority at this point.
@@ IX Conclusion
In conclusion, this paper introduces a novel baseline model, Valen, for solving visual abstract reasoning tasks, particularly Bongard-Logo problems and RPM problems. Valen, a probabilistic-highlighting model, demonstrates remarkable performance in handling both image clustering reasoning and image progression pattern reasoning. The paper further delves into the underlying mechanisms of probability-highlighting solvers, realizing that these solvers approximate solutions to reasoning problem instances as distributions. To bridge the discrepancies between the estimated solution distribution and the true correct solution distribution, the paper introduces three methods: Tine, Funny, and SBR.
  1. 1.
    The Tine algorithm conducts an in-depth analysis of the reasoning mechanism of probability-highlighting solvers and elucidates that the inadequate quality of auxiliary samples in reasoning problem instances often leads to the estimated distribution deviating from the correct solution distribution when using such end-to-end deep learning models to estimate the solution distribution for abstract reasoning problems. Consequently, the Tine method ingeniously incorporates the concept of adversarial learning into the solver’s training process, leveraging adversarial techniques to generate more judicious auxiliary samples, for the Valen model, thereby refining its solution distribution estimation process. Tine’s contributions lie not only in demonstrating the effectiveness and potential of adversarial learning frameworks in solver training, but also in revealing the crucial role of auxiliary samples in delineating the boundaries of probability-highlighted solver solution distributions, and highlighting the importance of the quality of negative examples for machine intelligence in learning abstract concepts.
  2. 2.
    The Funny method, embracing the spirit of the Tine algorithm, introduces additional reasoning image regression tasks for Valen to achieve mutual information supervision, thereby modeling the distribution of reasoning sample representations encoded by Valen as a Gaussian mixture distribution to enable it to more directly capture the true form of the correct solution distribution. The Half-Split Decoder designed in the Funny method has also made significant contributions to balancing regression and reasoning tasks, potentially opening up new opportunities for solving reasoning problems in the form of Image Generation Model. The Funny method emphasizes the importance of modeling the representation distribution of reasoning samples for enhancing the level of solver’s reasoning intelligence.
  3. 3.
    The SBR method further extends the idea of the Funny method. By modeling the distribution of Valen’s encoded progressive pattern representations as a more interpretable and orderly mixture of Gaussian distributions, the SBR method significantly enhances Valen’s reasoning accuracy and interpretability. The SBR method represents a novel attempt in the field of RPM solver training, embodying an innovative approach and mechanism to utilize metadata. It successfully reverses the previously observed phenomenon where directly introducing metadata as a supervisory signal to the solver actually impairs the reasoning process[25, 32, 37]. The SBR method emphasizes that for solvers dedicated to learning abstract concepts, it is crucial to meticulously plan the distribution of latent variables aimed at extracting these abstract concepts, as this significantly enhances their reasoning intelligence.
This paper demonstrates the importance of explicitly planning the distribution of solutions in solving visual abstract reasoning problems.